# SceneComplete: Open-World 3D Scene Completion in Complex Real World Environments for Robot Manipulation

### v1

@@ III Experiments
We additionally observe that the runtime of the current implementation of SceneComplete is about 30s per object; we expect this to improve with advances in models and GPUs.
There are no existing methods that solve the same problem, so we compare against an ablation of SceneComplete, which uses the same pipeline, but instead of using image-to-3D shape completion, we use the convex-hull of the partial point-cloud associated with each segmented image region $R_{i}$.
Figure 6: Qualitative results of grasp predictions on partial point clouds versus SceneComplete reconstructions. As can be seen, the parallel jaw grasps predicted using [13] that samples antipodal grasps on the SceneComplete reconstruction enables more robust predicts with less collisions on the ground truth (GT) scene. We also demonstrate dexterous grasps using both Shadow Hands and Allegro Hands [14], highlighting improved manipulation of complex objects with complete 3D reconstructions.
Figure 7: While impressive, these models [39] struggle with partial observability in real-world depth data even on in-category examples. An underlying assumption is that these models operate on partial inputs in category-specific canonical pose [40].
We believe that our results illustrate the ability of SceneComplete to:
  • produce highly accurate reconstructions of unobserved parts of a scene,
  • operate in cluttered scenes with substantial occlusion,
  • generalize “zero-shot” to previously unseen object types, and
  • provide important spatial information that supports collision-free manipulation and multi-fingered grasping.
III-A Scene Reconstruction
@@
We begin by evaluating the ability of SceneComplete to generate highly accurate scene reconstructions, using the GraspNet-1B dataset [12]. This dataset is particularly suitable for our evaluation due to its large collection of 190 cluttered tabletop scenes, featuring a diverse array of 88 unique objects in various configurations. It includes ground-truth 3D object models and poses for each scene, as well as real RGB-D images of the scenes.
Our primary reconstruction evaluation metric is based on comparing the ground-truth meshes with those produced by the SceneComplete. Let $U^{*}$ be the union of the volumes meshes of the ground-truth objects and let $\widehat{U}$ be the union of the volumes of the meshes produced by a reconstruction algorithm. Then the intersection-over-union metric is
$\text{IoU}(U^{*},\widehat{U})=\frac{U^{*}\cap\widehat{U}}{U^{*}\cup\widehat{U}}\;\;.$
We evaluated both SceneComplete and SceneComplete-NoComp (the ablation without 3D shape completion) on all 190 scenes from GraspNet-1Billion and found that SceneComplete had an average per-scene IoU score of 0.39 which is more robust, and substantially better than the SceneComplete-NoComp’s score of 0.14.
@@
In addition, we performed a basic test of the utility of this model for manipulation by using the antipodal grasping method of DexNet [13] to generate grasps on the objects in the reconstructed scene. To illustrate the importance of whole-scene reconstruction we then ask the question: of these potential grasps, which ones are in collision with other objects? Concretely, for a reconstructed scene, we (a) use DexNet to sample a set $G$ of potential grasps, (b) filter out all grasps in $G$ that would cause a collision in the reconstructed scene to obtain a subset $G^{\prime}$, (c) filter out all grasps in $G$ that would cause a collision in the ground truth scene to obtain a subset $G^{*}$, and compute $\lvert G^{\prime}-G^{*}\rvert/\lvert G^{\prime}\rvert$, which is the percentage of grasps that the reconstruction would allow, that in fact collide. We find that, on average over the 190 scenes, only 26% of grasps generated by SceneComplete are in collision with real objects, as opposed to 49% of those generated by SceneComplete-NoComp.
Fig. 6 shows a qualitative analysis of the same. As can be seen, there is a significant reduction in the number of collision of the grasps predicted on the reconstructed scene (SceneComplete) in the actual scene (GT scene) than grasps predicted on the partial scene.
III-B Dexterous Grasping and Stability
@@
A significant test of the utility of our approach is whether the individual object reconstructions support the computation of good dexterous grasps for a multi-fingered hand. We evaluate each object reconstruction as follows:
@@
  • Instantiate Isaac Gym [41] with the selected hand and the ground truth object mesh.
  • Lift the hand within the simulation and detect whether the object is dropped using PhysX as the physics engine.
We tabulate the percentage of such tests that succeed. We selected 20 objects arbitrarily from the Graspnet-1B dataset and evaluated SceneComplete against SceneComplete-NoComp on the ground-truth object shape (an important upper-bound that illustrates the reliability of DexGraspNet for selecting such grasps). We notice that the number of grasps sampled by DexGraspNet varies greatly based on the object geometry. Overall, we observe that DexGraspNet can sample 2×2$\times$2 × as many valid grasp proposals with SceneComplete than on SceneComplete-NoComp.
IV Discussion
@@
Although our results are very promising, there are of course many failure modes in a composition of so many modules, which can have cascading effects on overall system performance. We outline some failure modes and opportunities for improvement below. Prompting and segmentation: Occasionally the VLM fails to detect some of the object(s) in the image, causing the reconstruction of the object to be missed altogether. Tuning the prompt can mitigate this problem, as can prompting multiple ways for multiple hypotheses. Segmentation: Grounded-SAM occasionally segments parts of an object along with the full object which leads to multiple reconstruction hypotheses for the same object, or detects two cluttered objects together as one. We mitigate this partially using IoU-based de-duplication. Inpainting: Our current inpainting strategy operates on a relatively isolated object, which removes some important context. We mitigate this by slightly increasing the bounding box and adapting the model, but there is further room for improvement. Image-to-3D: Although remarkable, these models can sometimes fail to generate plausible reconstructions when given images from highly unusual viewpoints. Scaling and registration: Our scaling method is naive and would be improved in some cases by making it be non-isotropic. Registration sometimes fails on uniformly-textured objects, where it is difficult to find distinctive features. It can also fail if the scaling is incorrect.
Future work
One important strategy for making the system less error-prone is to move to a more generative setting with quantified uncertainty. If each module could generate multiple hypotheses, conditioned on its inputs, it would be possible to search for an interpretation that is collectively satisfactory to all the modules.
Conclusion
We have presented a system that solves full-scene reconstruction from a single real-world RBG-D input in cluttered, occluded scenes with no assumptions about object categories. We have built on an incredibly strong foundation of existing general-purpose open-domain perception models and believe our approach will be able to adapt to and profit from future advances in such models. One goal of this paper is to emphasize the importance of this problem for robot manipulation in real open-world environments and to encourage others to propose alternative solution strategies. We hope that overall advances in scene understanding in realistic manipulation settings will enable much more robust and capable robot manipulation systems.

### v4

@@ IV Experiments
We observe that the runtime of the current implementation of SceneComplete is about 20s per object on a single NVIDIA RTX 4090 GPU; we expect this to improve with advances in models and GPU architectures.
IV-A Scene Reconstruction and Grasping
We first evaluate SceneComplete on the task of reconstructing tabletop scenes from just a single RGB-D image.
@@ IV-A Scene Reconstruction and Grasping
Dataset: We use the GraspNet-1B dataset [1] for our evaluation. This dataset is particularly suitable for our setting, due to its large collection of 190 cluttered tabletop scenes, featuring 88 unique objects in various configurations. It includes ground-truth 3D object models and poses for each scene, as well as real RGB-D images.
Baseline: Our primary baselines for comparison are OctMAE [8] and ZeroGrasp [9]. OctMAE performs reconstruction by combining octree-based representation with a 3D Masked AutoEncoder (MAE). ZeroGrasp performs simultaneous 3D reconstruction and 6D grasp pose prediction using an octree-based CVAE. Both methods expect a single-view RGB-D image along with a corresponding foreground mask as input, and output the reconstructed scene. For our experiments, we provide OctMAE and ZeroGrasp with the ground-truth segmentation masks. We also report numbers against the partial point cloud (PartialDecomp).
Metrics: To measure 3D reconstruction quality and fidelity, we report well-known 3D metrics such as Chamfer distance (CD) and Earth Mover’s distance-maximum mean discrepancy (MMD-EMD) metric, similar to existing works [8, 19, 29]. Both CD and MMD-EMD metrics expect pointclouds as input. For SceneComplete, we sample points uniformly from the reconstructed object meshes. Since OctMAE predicts occupancy values, we use the reconstructed point cloud produced by its occupancy values, normal vectors, and SDF. We also report the Mesh Intersection-over-Union (MIoU) metric, which is based on comparing the ground-truth meshes with the reconstructed meshes. Specifically, let $U^{*}$ be the union of volumes enclosed by the ground truth object meshes, and let $\widehat{U}$ be the union of the volumes enclosed by the meshes produced by a reconstruction algorithm. Then the intersection-over-union metric between the meshes is
$\text{MIoU}(U^{*},\widehat{U})=\frac{U^{*}\cap\widehat{U}}{U^{*}\cup\widehat{U}}\;\;$.
MIoU explicitly penalizes both under-reconstructions (missing parts), over-reconstructions (excess geometry), and registration errors. To compute MIoU, we make the reconstructed meshes produced by SceneComplete, ZeroGrasp, and OctMAE watertight using ManifoldPlus [30].
@@
Collision-free Grasping: We perform a basic test of the utility of SceneComplete on an important downstream task of grasping, by using an antipodal grasp generation method [3] to generate grasps on the objects in the reconstructed scene. To illustrate the importance of whole-scene reconstruction on grasping, we ask the question: of these potential grasps, which ones are in collision with the ground-truth scene? Concretely, for a reconstructed scene, we (a) sample a set $G$ of collision-free grasps using antipodal sampling, (b) evaluate grasps in $G$ that would cause a collision in the ground truth scene to obtain a subset $G^{\prime}$, and compute the Grasp Collision metric GC as ($\lvert G^{\prime}\rvert/\lvert G\rvert$) which is the percentage of grasps that the reconstruction would allow, that in fact collide, similar to the metric adopted by other works [31, 19]. In our case, we set $G$ to 40.
Reconstruction Grasping
[…]
PartialDecomp 0.166 3.16 3.32 53.5
OctMAE [8] 0.445 1.73 3.11 20.3
ZeroGrasp [9] 0.440 1.86 3.07 18.9
SceneComplete 0.478 1.54 3.06 16.4
TABLE I: Comparison of shape reconstruction methods on GraspNet-1B. Higher MIoU indicates better shape fidelity. Lower CD, MMD-EMD, and GC indicate more accurate and feasible reconstructions for downstream grasping. CD and MMD-EMD are scaled by $10^{4}$ and $10^{2}$ respectively.
[…]
PartialDecomp $0.46\pm 0.34$ $0.17\pm 0.13$ 0.32
SceneComplete $\textbf{0.81}\pm\textbf{0.2}$ $\textbf{0.73}\pm\textbf{0.18}$ 0.77
[…]
Results: As shown in Table I, our method outperforms the baseline methods in all metrics. We also visualize the comparisons in Figure 5. While OctMAE produces visually plausible reconstructions, it sometimes fails to recover parts of objects that are occluded by other objects, or are partially observable due to the viewpoint. This also results in higher grasp collisions as shown in Table I, while our method recovers such missing regions. Moreover, since OctMAE directly predicts a scene-level reconstruction, it often hallucinates geometry connecting distinct objects (highlighted in Fig. 5), which results in a lower MIoU. In contrast, our object-centric approach reconstructs each object individually, preserving a clear separation between them. We find that on average, only 16% of grasps generated by SceneComplete are in collision with real objects, as opposed to 20% of those generated by OctMAE, 19% of those generated by ZeroGrasp, and 53% of those generated by PartialDecomp.
Figure 8: Qualitative comparisons of scene reconstructions on the scans collected in our lab. For each scene we show, the input RGB-D image, OctMAE reconstructions, our reconstructions, grasp proposals on input partial point cloud, grasp proposals on OctMAE’s reconstructions, and grasp proposals on our reconstructions. We show the scene from a front and back viewpoint.
IV-B Object Grasping and Dexterous Manipulation
We next evaluate the utility of SceneComplete for object grasping and dexterous manipulation in simulation.
Dataset: We use a subset of the YCB-Video [2] dataset consisting of 30 cluttered tabletop scenes with 4 to 5 objects per scene from the YCB [17] dataset for simulated grasping experiments using a parallel jaw gripper. For dexterous manipulation, we evaluate on 20 object instances chosen arbitrarily from the GraspNet-1B [1] dataset.
Baseline: We compare against the raw input partial point cloud (PartialDecomp) as the baseline. Since this experiment is object-centric, its not compatible with OctMAE, which produces scene-level reconstruction without per-object separation. While instance segmentation could in principle be used to individuate objects, they do not reliably recover the extent or shape of each reconstructed object, making fair comparison infeasible for object-centric manipulation tasks. We show comparisons with OctMAE on a scene-level grasping task in the next section.
Metrics: To assess the effectiveness of SceneComplete for object manipulation, we evaluate the Grasp Success Rate (GSR) using simulated grasp attempts (a common metric adopted by prior works [32, 33]) in Isaac Gym. Our scene consists of a Franka Emika Panda arm with a parallel jaw gripper, and scenes from the YCB-V dataset [2]. We compare the GSR achieved by SceneComplete against PartialDecomp and adopt two established and distinct grasping methods for generating grasp proposals:
  • Antipodal grasping using [3]: This method samples antipodal points to generate collision-free grasps on the object meshes. To make the objects compatible with simulation and to improve efficiency, we perform an approximate convex decomposition of the watertight object meshes into convex hulls using CoACD [34].
  • Contact-GraspNet [4]: This method generates grasp proposals directly from point clouds and was trained to predict grasps from partial observations. We generate grasp proposals on both the reconstructed point cloud (from SceneComplete) and the raw partial point cloud, and evaluate these grasps on the ground-truth scenes.
For each grasping method, we first generate candidate grasps on both SceneComplete’s reconstructed scene and PartialDecomp. For each grasp, we simulate a pick attempt on the reconstructed meshes in Isaac Gym and retain only those grasps for which the object remains securely held in the gripper. From these filtered grasps, we randomly select upto $40$ grasps per object and evaluate them on the ground-truth object meshes in Isaac Gym. Each evaluation consists of successfully picking up the object from its computed grasp and holding it in the air without dropping, as shown in Fig. 6. GSR is computed as the proportion of successful grasp attempts out of the $40$ evaluated grasps. Objects that do not yield valid grasps due to constraints such as exceeding the gripper width are excluded from our evaluation.
@@ IV-B Object Grasping and Dexterous Manipulation
Dexterous Grasping and Stability: A significant test of the utility of our approach is whether the object reconstructions support the computation of good dexterous grasps for a multi-fingered hand. We evaluate object reconstruction as:
@@
  • Instantiate Isaac Gym with the selected hand and the ground truth object mesh.
  • Similar to other methods that evaluate dexterous grasping [5, 35], we lift the hand and rotate it within the simulation, and detect whether the object is dropped using PhysX as the physics engine. We visualize dexterous grasps on representative objects using both Shadow and Allegro hands in Fig. 7.
We calculate the percentage of such tests that succeed. For evaluation, we selected 20 objects from the GraspNet-1B dataset and evaluated SceneComplete against PartialDecomp on ground-truth objects—an important upper bound illustrating the reliability of DexGraspNet for selecting such grasps.
Results: As shown in Table II, reconstructing scenes with SceneComplete significantly improves grasp success rates in simulation using a parallel jaw gripper. Across both grasping methods, SceneComplete achieves over twice the number of successful grasps compared to those from partial input alone. For dexterous grasping, we observe that the number of stable grasps sampled by DexGraspNet varies with object geometry. On average, we evaluate up to 100 random grasps per object and find that SceneComplete enables twice as many valid dexterous grasps compared to PartialDecomp. These improvements show that reconstructing scenes enables more reliable grasping and manipulation in cluttered environments.
IV-C Real Robot Experiments
We validate the real-world applicability of SceneComplete through experiments on a physical robot. Our experimental setup includes a Franka Emika Panda arm equipped with a wrist-mounted Intel RealSense D435 camera.
Dataset: We evaluate our method on 15 distinct tabletop scenes collected in our lab, each containing 4 to 6 everyday objects, specifically chosen to minimize their likelihood of appearing in the training distribution of the methods.
Baseline: We compare against PartialDecomp and OctMAE in our evaluation.
Metrics: For each scene, we capture an initial RGB-D image using the wrist-mounted camera and randomly select and execute kinematically feasible and collision-free grasps on each object in the scene. We measure the success rate as the percentage of objects that were successfully picked up and report the results in Table III.
Results: Our method achieves a success rate of $\sim$73%, significantly outperforming OctMAE and grasps generated directly from the input partial point cloud alone. We show qualitative comparisons with OctMAE in Fig. 8 and additional results in the supplementary. OctMAE struggles to produce plausible reconstructions on these everyday objects, often hallucinating geometry between distinct objects, leading to no valid grasps being generated for some objects. SceneComplete on the other hand, produces object reconstructions that align closely with the ground truth scene. We note that for SceneComplete, most grasping failures occur due to inaccuracies in the estimated object size, which subsequently leads to errors in registration. In general, SceneComplete allows robust manipulation of objects in cluttered real-world settings.
Partial OctMAE [8] SceneComplete
[…]
TABLE III: Real robot success rate (%) measured as the percentage of objects picked successfully by the real robot for each method.
V Discussion
@@
Although our results are very promising, there are of course many failure modes in a composition of so many modules, which can have cascading effects on overall system performance. We outline some failure modes and opportunities for improvement. Prompting and segmentation: Occasionally the VLM fails to detect some of the object(s) in the image. Tuning the prompt mitigates the problem, as can prompting multiple ways for multiple hypotheses. Segmentation: Grounded-SAM occasionally segments parts of an object along with the full object which leads to multiple reconstruction hypotheses for the same object. We mitigate this partially using IoU-based de-duplication. Inpainting: Our current inpainting strategy operates on a relatively isolated object, which removes some important context. We mitigate this by slightly increasing the bounding box and adapting the model, but there is room for improvement. Image-to-3D: Although remarkable, these models can sometimes fail to generate plausible reconstructions when given images from highly unusual viewpoints. Scaling and registration: Our scaling method is naive and would be improved in some cases by making it non-isotropic. Registration sometimes fails on uniformly-textured objects, where it is difficult to find distinctive features. Seed value: The performance of the image-to-3D and inpainting models may sometimes vary depending on the seed value selected.
Conclusion & Future Work
We have presented a system that solves full-scene reconstruction from a single real-world RBG-D input in cluttered, occluded scenes with no assumptions about object categories. We have built on an incredibly strong foundation of existing general-purpose open-domain perception models and believe our approach will be able to adapt to and profit from future advances in such models. One goal of this paper is to emphasize the importance of this problem for robot manipulation in real open-world environments and to encourage others to propose alternative solution strategies. We hope that overall advances in scene understanding in realistic manipulation settings will enable much more robust and capable robot manipulation systems. One important strategy for making the system less error-prone is to move to a more generative setting with quantified uncertainty. If each module could generate multiple hypotheses, conditioned on its inputs, it would be possible to search for an interpretation that is collectively high-probability for all the modules.
VI Acknowledgments
We gratefully acknowledge support from NSF grant 2214177; from AFOSR grant FA9550-22-1-0249; from ONR MURI grants N00014-22-1-2740 and N00014-24-1-2603; from the MIT Quest for Intelligence; and from the Robotics and Artificial Intelligence Institute.